OpenStudy (anonymous):

from a base elevation of 6800 feet, a mountain peak in colorado rises to a summit elevation of 14,260 feet over a horizontal distance of 15,847 feet. Find the grad of the peak

OpenStudy (anonymous):

\[\Delta y/\Delta x=(14260-6800)/(15847)=.471\]

OpenStudy (anonymous):

that is the rate the elevation is changing with respect to horizontal distance

OpenStudy (anonymous):

Thank you, you said it's the rate, but is that the same as the grade

OpenStudy (anonymous):

find the slope if is exists to the pair containing the pair of points (-4,-10) and (-5,-11)

OpenStudy (anonymous):

(-11-(-10))/(-5-(-4))=-1/(-1)=1